OpenStudy (anonymous):

how to find the slope intercept form of y=-2x+2

OpenStudy (agreene):

it is in slope intercept form: y=mx+b with: m is slope b is y intercept.

OpenStudy (anonymous):

so that means slope is -2 and y intercept is 2

OpenStudy (agreene):

yes.

OpenStudy (anonymous):

So as agreene said.... m is ur slope, so ur m in this equation is -2

OpenStudy (anonymous):

thanks so it would be graphed as (-2,2)

OpenStudy (anonymous):

So, ur b is the starting point so it would be (2,-2)

OpenStudy (anonymous):

b = y intercept

OpenStudy (anonymous):

yea but dont i still have to have another point on my graph because it says i have to have a straight line so what would the other point be?

OpenStudy (anonymous):

sorry if im confusing

OpenStudy (anonymous):

can I help?

OpenStudy (anonymous):

do you use a ti-84 calculator?

OpenStudy (anonymous):

yes you can help i really need it and no sorry :/

OpenStudy (anonymous):

I could still help give me a second

OpenStudy (anonymous):

okay thanks, i'll wait...

OpenStudy (anonymous):

formula for slope intercept form is y=mx+b familiar with this?

OpenStudy (anonymous):

so do you also need y intercept?

OpenStudy (anonymous):

yes im familiar with it and yes i also need the y intercept please

OpenStudy (anonymous):

for y intercept is y=-2(0)+2 since you are solving for y the x value should be zero.

OpenStudy (anonymous):

so the answer that i graph should be (-2,2)

OpenStudy (anonymous):

that's if you need to solve for both x and y intercept. But it's better to see it by step so you could understand it better.

OpenStudy (anonymous):

no i only need to solve for y intercep so the answer would be?

OpenStudy (anonymous):

So I will continue

OpenStudy (anonymous):

okay thanks..

OpenStudy (anonymous):

if only y intercept then is (0,2)

OpenStudy (anonymous):

the slope is -2

OpenStudy (anonymous):

so i would have to graph it as (0,2) then rise -2 ?

OpenStudy (anonymous):

m=-2 y intercept is (0,2)

OpenStudy (anonymous):

yes but do you know how to graph the slope

OpenStudy (anonymous):

I'll draw the graph

OpenStudy (anonymous):

so that means i graph (0,2) then run over -2 ?

OpenStudy (anonymous):



OpenStudy (anonymous):

okay i graphed the 0,2 but how do igraph the -2